OpenStudy (anonymous):

what is 1/8 plus 1/9? with work

OpenStudy (anonymous):

first you will need to find the Least Common Denominator.

OpenStudy (anonymous):

which i think is 72 so you will need to multiply \[\frac{ 1 }{ 8 }*\frac{ 9 }{ 9}\] and \[\frac{ 1 }{ 9 }*\frac{ 8 }{ 8}\]

OpenStudy (anonymous):

and you will get\[\frac{ 9 }{ 72}+\frac{ 8 }{ 72 }=\frac{ 17 }{ 72 }\]

OpenStudy (anonymous):

hope this somewhat helps